ADNAN MUHAMMED

Grade 12,

a boy whirls the stone in horizontal circle of radius 1.5m....at a height of 2 m,,,,,above the ground..the string breaks and the stone flies off horizontally and strikes the ground after travellin a hori distance of 10m,,,,find its centripeal acceleration..

a boy whirls the stone in horizontal circle of radius 1.5m....at a height of 2 m,,,,,above the ground..the string breaks and the stone flies off horizontally and strikes the ground after travellin a hori distance of 10m,,,,find its centripeal acceleration..

Grade:11

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear shivanshi

It is a question of projectile motion

let the horizontal velocity of stone when string breaks is =V

verticale velocity is zero

let  time taken to reach the ground is t

equation for verticle motion

 h=1/2gt2

2=1/2 * 10 * t2

t= √.4

so equation for horizontal motion

S =Vt

10 = V *√.4

V= 15.81 m/sec

so centripital acceleration = V2/r

                                      = (15.81)2 /1.5

                                      =166.64 m/sec2

Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very

quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin